OpenStudy (anonymous):

2^-9x=7^x+10 Solve for x and write the exact answer using base-10 logarithms.

OpenStudy (anonymous):

\[2^{-9x}=7^{x+10}\]

OpenStudy (anonymous):

Yeah that's how it is written.

OpenStudy (anonymous):

I got some weird funky fraction and I don't think it's right.

OpenStudy (anonymous):

Task steps: Take the log of both sides. Use the power rule to pull the variables out of the logarithm. Distribute where possible. Move terms with x to one side of the equation, other terms to the other side. Factor out x. Divide by the other factor to get x=.... It will probably be a mess of logs and numbers, but that's the answer. If you need an approximation, use your calculator.

OpenStudy (anonymous):

What answer did you get?

OpenStudy (anonymous):

Does the answer have the term "ln" in it?

OpenStudy (anonymous):

10ln(7)/9ln(2)+ln(7)

OpenStudy (anonymous):

\[2^{-9x}=7^{x+10} \implies -9x \log 2=(x+10)\log 7 \implies xlog7+9xlog2=-10\log7\]\[\implies x(\log7+9\log2)=-10\log7 \implies x=\frac{-10\log7}{\log7+9\log2}\]

OpenStudy (anonymous):

so ln is another term for log, okay thank you!

OpenStudy (anonymous):

The specification in the problem said to use common log. I think your answer is within a negative sign of the correct answer.

OpenStudy (anonymous):

ln is the natural log (base e). The common log is base 10.

OpenStudy (anonymous):

Oh okay, thanks for your help!

OpenStudy (anonymous):

No sweat. Do math every day.